# Training Set Reconstruction from Differentially Private Forests: How Effective is DP?

### v2

@@ Abstract
Recent research has shown that machine learning models are vulnerable to privacy attacks targeting their training data. To mitigate these risks, differential privacy (DP) has become a widely adopted countermeasure, as it offers rigorous privacy protection. In this paper, we introduce a reconstruction attack targeting state-of-the-art $\varepsilon$-DP random forests. By leveraging a constraint programming model that incorporates knowledge of the forest’s structure and DP mechanism characteristics, our approach formally reconstructs the most likely dataset that could have produced a given forest. Through extensive computational experiments, we examine the interplay between model utility, privacy guarantees and reconstruction accuracy across various configurations. Our results reveal that random forests trained with meaningful DP guarantees can still leak portions of their training data. Specifically, while DP reduces the success of reconstruction attacks, the only forests fully robust to our attack exhibit predictive performance no better than a constant classifier. Building on these insights, we provide practical recommendations for the construction of DP random forests that are more resilient to reconstruction attacks and maintain non-trivial predictive performance.
1 Introduction
@@
The success of machine learning algorithms in critical applications, such as medicine, kidney attribution and credit scoring [Dildar et al., 2021, Aziz et al., 2021, Dastile et al., 2020], hinges on access to large volumes of sensitive data, including medical records, personal histories and financial information. This extensive data is crucial for training models capable of achieving high accuracy. At the same time, trustworthiness in these applications demands transparency, as emphasized by regulations like Article 13 of the EU AI Act [^1] https://artificialintelligenceact.eu/article/13/. Transparency involves the ability to audit models and explain their outputs and inner workings, which fosters public confidence and ensures compliance with legal requirements.
@@
However, transparency can inadvertently create vulnerabilities. When models are made public, even via black-box APIs, they become susceptible to privacy attacks that can infer sensitive information about their training data [Rigaki and Garcia, 2023]. Recent studies on reconstruction attacks [Dwork et al., 2017] demonstrated that large language models [Carlini et al., 2021], diffusion models [Carlini et al., 2023], neural networks [Haim et al., 2022], random forests (RFs) and a wide range of interpretable models [Ferry et al., 2024a, b] inherently leak part of their training data through their structure or outputs.
@@
To mitigate such privacy risks, differential privacy (DP) [Dwork and Roth, 2014] has emerged as the de facto standard, as it limits the amount of information an adversary can gain regarding any single individual in the dataset. DP theoretically bounds the success of membership inference attacks (which aim at inferring the absence or presence of a given individual in a model’s training data), but also empirically provides a form of protection against other attacks. For instance, recent work on deep learning for medical imaging [Ziller et al., 2024] has shown that DP mitigates the success of some reconstruction attacks, even for loose privacy guarantees. However, DP imposes significant trade-offs as ensuring privacy involves adding noise or other modifications that degrade a model’s predictive performance. Consequently, practitioners must carefully balance privacy guarantees and model utility in sensitive domains.
@@
In this paper, we focus on reconstruction attacks targeting tree ensembles, such as RFs. These models frequently achieve state-of-the-art performance on tabular data, but recent research [Ferry et al., 2024b] has revealed that their structure can inherently expose their training data. While several DP tree-based models have been designed [Fletcher and Islam, 2019], it remains unclear how well DP can effectively prevent such privacy leaks. To address this gap, the main contributions of our study are:
@@
  • We introduce an effective reconstruction attack against state-of-the-art DP RFs. Building on the attack paradigm pioneered by [Ferry et al., 2024b], our approach frames the reconstruction attack as a Constraint Programming (CP) model, which identifies a most likely dataset capable of generating the observed RF and noisy counts.
@@
  • Through extensive experiments, we show that DP RFs, while formally satisfying privacy guarantees, can still be exploited to retrieve portions of their training data. In particular, all investigated DP RFs with non-trivial predictive performance reveal information about their training data, which often extends beyond distributional patterns to specificities of the dataset used. To enable the reproducibility of all the experiments, the source code of our method is openly available at https://github.com/vidalt/DRAFT-DP, under a MIT license.
@@
2 Technical Background
2.1 Differential Privacy
Originally formalized in Dwork [2006], DP enables the computation of useful statistics on private data while strictly limiting the information that can be inferred about any single individual. A randomized mechanism $\mathcal{K}$ satisfies $\varepsilon$-DP if, for any pair of datasets $D_{1}$ and $D_{2}$ differing by at most one record, and all $U\subseteq\text{Range}(\mathcal{K})$:
@@ 2.1 Differential Privacy
Let $q$ represent a deterministic query such as counts, sums, averages, or, in the context of tree ensembles, the number of samples of a particular class in a data subset. DP can give protected access to this query through a probabilistic mechanism $\mathcal{K}$, which outputs $q$ with additional noise. As detailed in Dwork et al. [2016], the amount of noise needed to protect the data depends on the global sensitivity of the query, defined as $S(q)=\max||q(D_{1})-q(D_{2})||_{1},$ in which the maximum is taken over all pairs of datasets $D_{1}$ and $D_{2}$ differing in at most one element. Thus, $S(q)$ quantifies the largest possible change in the query output due to a single record’s addition or removal (e.g., $S(q)=1$ for counts). For instance, the Laplace mechanism, a widely used DP method, adds noise sampled from a Laplace distribution:
@@
in which $Y_{i}$ are random variables i.i.d. following the Laplace distribution $\text{Lap}(S(q)/\varepsilon)$ with $\varepsilon$ the privacy budget granted to the mechanism.
When multiple DP mechanisms are combined, the overall privacy budget is governed by two fundamental composition theorems [McSherry and Talwar, 2007]. First, sequential composition states that applying several $\varepsilon_{i}$-DP mechanisms to the same dataset results in $\sum_{i}\varepsilon_{i}$-DP. In contrast, parallel composition states that applying several $\varepsilon_{i}$-DP mechanisms to disjoint subsets of data results in $\max_{i}\varepsilon_{i}$-DP. Finally, the postprocessing property ensures that any function applied to the output of $\mathcal{K}$ does not affect the DP guarantee, providing flexibility for downstream analysis.
DP mechanisms have been widely adopted to protect privacy when releasing machine learning models. For example, noise can be added to gradients during neural network training [Abadi et al., 2016], or split thresholds can be perturbed in decision tree predictors [Fletcher and Islam, 2019]. In particular, DP can mitigate privacy risks such as membership inference or model inversion [Ye et al., 2022], and implementations of DP mechanisms have been integrated within popular libraries, such as IBM’s diffprivlib [Holohan et al., 2019] or Google’s differential-privacy [Google, ].
2.2 Differentially-Private Random Forests
@@
In most random forest implementations, including popular libraries such as scikit-learn [Pedregosa et al., 2011], predictions are made using soft voting: each tree outputs class probabilities based on the normalized class counts in its leaves and the ensemble prediction averages these probabilities. Specifically, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ in tree $t$, the model stores $n_{tvc}$, the number of training examples of class $c$ that reach $v$. This information is essential for on-device inference and audits but also introduces risks, as it exposes model parameters to potential adversaries [Ferry et al., 2024b].
To address these risks, an extensive body of research has focused on constructing DP RFs to conciliate model accessibility and privacy protection, with many of these seminal methods being surveyed by Fletcher and Islam [2019]. Among these, the method proposed by Jagannathan et al. [2012], which combines randomized tree structure and noisy counts publication, was the first to achieve strong empirical predictive performance and formal DP guarantees. Consequently, we use it as a foundation of our experimental study. In this approach, each tree is trained on the complete dataset $D$. Each tree of a chosen depth $d$ is grown by randomly selecting attributes and split values for all internal nodes $v\in\smash{\mathcal{V}_{t}^{I}}$, avoiding direct access to $D$ during this phase to preserve the DP budget. Subsequently, the class counts $n_{tvc}$ at each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are obfuscated using the Laplace mechanism [Dwork et al., 2016] and cast to integer values, as illustrated in Figure 1.
@@
[…]
3 Reconstruction Attack against DP RFs
@@
Attack Model. We assume white-box access to an $\varepsilon$-DP RF constructed using the algorithm of Jagannathan et al. [2012], as described in Section 2.2. This includes access to the structure of the forest (split attributes and values) and the (noisy) counts at the leaves. Consistent with the literature on reconstruction attacks [e.g., Dwork et al., 2017, Ferry et al., 2024b], we also assume knowledge of the number of attributes $M$ and training examples $N$. While the knowledge of $M$ is necessary for inference, the knowledge of $N$ simplifies the attack but is not strictly required. In particular, in Appendix A, we demonstrate that our attack can reconstruct the dataset without prior knowledge of $N$, making it the first method to target full dataset reconstruction without this assumption. We also assume the privacy budget $\varepsilon$ is known. This does not degrade the DP guarantees, is generally disclosed by practitioners [^2] For example, Desfontaines [2021] documents real-world applications of DP and their associated budgets., and researchers have advocated this for transparency [Dwork et al., 2019].
@@
Reconstruction Approach. Our approach is grounded on the CP formulation introduced in Ferry et al. [2024b], with additional constraints, variables, and an objective specifically designed to model the (observed) noisy counts, the underlying (guessed) true counts, and their likelihood. Solving the original problem formulation (called DRAFT) was shown to be NP-Hard. Our reconstruction problem for DP RFs can be seen as a generalization (i.e., DRAFT is a special case with no noise) and is also NP-hard, though scalable with state-of-the-art CP solvers [Perron and Didier, 2024].
@@
Our formulation revolves around three main groups of decision variables. The objective of a CP solver is to find compatible variable values that satisfy a given set of constraints, yielding a feasible solution. Moreover, as explained later in this section, an objective function steering the search towards likely solutions (reconstructions) will be included. First, for each tree $t\in\mathcal{T}$, the (guessed) true number of examples of class $c\in\mathcal{\mathcal{C}}$ in each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are encoded using a variable $n_{tvc}\in\mathbb{N}$. Next, a set of variables encodes each reconstructed example $k\in\{1,\dots,N\}$: $x_{ki}\in\{0,1\}$ encodes the value of its $i^{\text{th}}$ attribute $i\in\{1,\dots,M\}$, while $z_{kc}\in\{0,1\}$ indicates if it belongs to class $c\in\mathcal{C}$. Third, the path of each example $k\in\{1,\dots,N\}$ through each tree $t\in\mathcal{T}$ is modelled using a binary variable $y_{tvkc}\in\{0,1\}$, which is $1$ if it is classified by leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ as class $c\in\mathcal{C}$, and 00 otherwise. The following constraints linking these variables are then enforced: $\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}n_{tvc}=N$ $t\in\mathcal{T}$ (1) $\sum\limits_{c\in\mathcal{C}}z_{kc}=1$ $k\in\{1,\dots,N\}$ (2) $z_{kc}=\sum\limits_{\smash{\mathclap{v\in\smash{\mathcal{V}_{t}^{L}}}}}y_{tvkc}$ $k\in\{1,\dots,N\},t\in\mathcal{T},c\in\mathcal{C}$ (3) $\sum_{k=1}^{\smash{N}}y_{tvkc}=n_{tvc}$ $t\in\mathcal{T},~{}v\in\smash{\mathcal{V}_{t}^{L}},~{}c\in\mathcal{C}$ (4)
@@
In a nutshell, Constraint (1) ensures that the example counts sum up to $N$ within each tree. Constraint (2) guarantees that each example is associated to exactly one class. Constraint (3) enforces that each example contributes only to the counts of its assigned class. Moreover, Constraint (4) ensures consistency between the counts at each leaf and the examples assigned to it. Finally, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ of each tree $t\in\mathcal{T}$, let $\Phi_{v}^{+}$ (respectively $\Phi_{v}^{-}$) be the set of binary attributes that must equal $1$ (respectively 00) for an example to fall into $v$. Then, the following constraint enforces consistency between the examples’ assignments to the leaves, and the value of their attributes:
@@
$\sum\limits_{c\in\mathcal{C}}y_{tvkc}=1\Rightarrow$ $\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{+}}x_{ki}=1\right)\wedge\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{-}}x_{ki}=0\right)$ $t\in\mathcal{T},~{}v\in\smash{\mathcal{V}_{t}^{L}},~{}k\in\{1,\dots,N\}$ (5)
@@
[…]
In practice, we pre-compute constants $p_{l}$ for all values $l\in\{-\delta,\dots,\delta\}$ such that $\mathbb{P}(\textrm{int}(Y_{tvc})\in\{-\delta,\dots,\delta\})\geq 0.999$. Details on the computation of $\delta$ are provided in Appendix B, but typically, $\delta=\lceil 12/\varepsilon_{v}\rceil$ is sufficient. With this, we can additionally restrict the feasible domain of the variables to $n_{tvc}\in\{\max(0,n_{tvc}^{*}-\delta),\dots,n_{tvc}^{*}+\delta\}$ and $\Delta_{tvc}\in\{-\delta,\dots,\delta\}$ and formulate the objective as:
$\max\ \sum\limits_{t\in\mathcal{T}}\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}\sum\limits_{l=-\delta}^{\delta}\log(p_{l})\mathds{1}_{\Delta_{tvc}=l},$ (9)
@@
4 Experimental Study

### v4

@@ Abstract
Recent research has shown that structured machine learning models such as tree ensembles are vulnerable to privacy attacks targeting their training data. To mitigate these risks, differential privacy (DP) has become a widely adopted countermeasure, as it offers rigorous privacy protection. In this paper, we introduce a reconstruction attack targeting state-of-the-art $\varepsilon$-DP random forests. By leveraging a constraint programming model that incorporates knowledge of the forest’s structure and DP mechanism characteristics, our approach formally reconstructs the most likely dataset that could have produced a given forest. Through extensive computational experiments, we examine the interplay between model utility, privacy guarantees and reconstruction accuracy across various configurations. Our results reveal that random forests trained with meaningful DP guarantees can still leak portions of their training data. Specifically, while DP reduces the success of reconstruction attacks, the only forests fully robust to our attack exhibit predictive performance no better than a constant classifier. Building on these insights, we also provide practical recommendations for the construction of DP random forests that are more resilient to reconstruction attacks while maintaining a non-trivial predictive performance.
I Introduction
@@
The success of machine learning algorithms in critical applications, such as medicine, kidney attribution and credit scoring [1, 2, 3], hinges on access to large volumes of sensitive data, including medical records, personal histories and financial information. This extensive data is crucial for training models capable of achieving high accuracy. At the same time, trustworthiness in these applications demands transparency, as emphasized by regulations like Article 13 of the EU AI Act. [^1] https://artificialintelligenceact.eu/article/13/ Transparency involves the ability to audit models and explain their outputs and inner workings, which fosters public confidence and ensures compliance with legal requirements.
@@
However, transparency can inadvertently create vulnerabilities by exposing new attack surfaces. When models are made public, even via black-box APIs, they become susceptible to privacy attacks that can infer sensitive information about their training data [4]. Recent studies on reconstruction attacks [5] demonstrated that large language models [6], diffusion models [7], neural networks [8], random forests (RFs) [9] and a wide range of interpretable models [10] inherently leak part of their training data through their structure or outputs.
@@
To mitigate such privacy risks, differential privacy (DP) [11] has emerged as the de facto standard, as it limits the amount of information an adversary can gain regarding any single individual in the dataset. DP theoretically bounds the success of membership inference attacks (which aim at inferring the absence or presence of a given individual in a model’s training data), but also empirically provides a form of protection against other attacks. For instance, recent work on deep learning for medical imaging [12] has shown that DP mitigates the success of some reconstruction attacks, even for loose privacy guarantees. However, DP imposes significant trade-offs as ensuring privacy involves adding noise or other modifications that degrade a model’s predictive performance. Consequently, practitioners must carefully balance privacy guarantees and model utility in sensitive domains.
@@
In this paper, we focus on reconstruction attacks targeting tree ensembles, such as RFs. These models frequently achieve state-of-the-art performance on tabular data, as evidenced in recent works [13] and competitions [14]. Furthermore, they are widely used in real-world deployments as they offer competitive storage costs and inference times. Their structured nature makes them particularly suitable to model audits, further facilitating their use in sensitive contexts. However, recent research [9] has revealed that their structure can also inherently expose their training data. While several DP tree-based models have been designed [15], it remains unclear how well DP can effectively prevent such privacy leaks.
To address this gap, our setting is designed to investigate the following fundamental question: under what conditions can a trained model be shared in a white-box manner without compromising the privacy of its training data? By considering state-of-the-art machine learning models for tabular data (RFs), alongside standard privacy protection mechanisms (pure $\varepsilon$-DP using the Laplace mechanism and basic composition), our results establish meaningful milestones for addressing this question. More precisely, our main contributions are:
@@
  • We introduce an effective reconstruction attack against state-of-the-art DP RFs. Building on the attack paradigm pioneered by [9], our approach frames the reconstruction attack as a Constraint Programming (CP) model, which identifies a most likely dataset capable of generating the observed RF and noisy counts.
@@
  • Through extensive experiments, we show that DP RFs satisfying pure $\varepsilon$-DP guarantees can still be exploited to retrieve portions of their training data. In particular, all investigated DP RFs with non-trivial predictive performance reveal information about their training data, often extending beyond distributional patterns to specificities of the dataset such as complete training examples. The source code of our method is available at https://github.com/vidalt/DRAFT-DP under an MIT license.
@@
The remainder of the paper is organized as follows. First, in Section II, we provide the necessary background on DP and describe the considered DP RFs algorithm. Then, we introduce our reconstruction attack model and formulation in Section III. Afterwards, we empirically evaluate its effectiveness in Section IV. Finally, we review the relevant related works on reconstruction attacks, post-processing of DP values and DP RFs in Section V before concluding and discussing promising research avenues in Section VI.
II Background Notions
II-A Differential Privacy
Originally formalized in [16], DP enables the computation of useful statistics on private data while strictly limiting the information that can be inferred about any single individual. A randomized mechanism $\mathcal{K}$ satisfies $\varepsilon$-DP if, for any pair of datasets $D_{1}$ and $D_{2}$ differing by the addition or removal of at most one record, and all $U\subseteq\text{Range}(\mathcal{K})$:
@@ II-A Differential Privacy
[…]
@@
in which $Y_{i}$ are random variables i.i.d. following the Laplace distribution $\text{Lap}(S(q)/\varepsilon)$ with $\varepsilon$ the privacy budget granted to the mechanism. When multiple DP mechanisms are combined, the overall privacy budget is governed by two fundamental composition theorems [18]. First, sequential composition states that applying several $\varepsilon_{i}$-DP mechanisms to the same dataset results in $\sum_{i}\varepsilon_{i}$-DP. In contrast, parallel composition states that applying several $\varepsilon_{i}$-DP mechanisms to disjoint subsets of data results in $\max_{i}\varepsilon_{i}$-DP. Finally, the postprocessing property ensures that any function applied to the output of $\mathcal{K}$ does not affect the DP guarantee, providing flexibility for downstream analysis.
DP mechanisms have been widely adopted to protect privacy when releasing machine learning models. For example, noise can be added to gradients during neural network training [19], or split thresholds can be perturbed in decision tree predictors [15]. In particular, DP can mitigate privacy risks such as membership inference or model inversion [20], and implementations of DP mechanisms have been integrated within popular libraries, such as IBM’s diffprivlib [21] or Google’s differential-privacy [22].
II-B Differentially-Private Random Forests
@@
In most random forest implementations, including popular libraries such as scikit-learn [23], predictions are made using soft voting: each tree outputs class probabilities based on the normalized class counts in its leaves and the ensemble prediction averages these probabilities. Specifically, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ in tree $t$, the model stores $n_{tvc}$, the number of training examples of class $c$ that reach $v$. This information is essential for on-device inferences and audits but also introduces risks, as it exposes model parameters to potential adversaries [9].
To address these risks, an extensive body of research has focused on constructing DP RFs to conciliate model accessibility and privacy protection, with many of these seminal methods being surveyed by [15]. Among these, the method proposed by [24], which combines randomized tree structure and noisy counts publication, was the first to achieve strong empirical predictive performance and formal DP guarantees. Consequently, we use it as a foundation of our experimental study. In this approach, each tree is trained on the complete dataset $D$. Each tree of a chosen depth $d$ is grown by randomly selecting attributes and split values for all internal nodes $v\in\smash{\mathcal{V}_{t}^{I}}$, avoiding direct access to $D$ during this phase to preserve the DP budget. Subsequently, the class counts $n_{tvc}$ at each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are computed using $D$. Finally, they are obfuscated using the Laplace mechanism [17] and cast to integer values, as illustrated in Figure 1.
@@
in which $Y_{tvc}$ is a random variable sampled from a Laplace distribution $\text{Lap}(1/\varepsilon_{v})$ and $\textrm{int}(.)$ is the integer part function. Appendix C-B provides qualitative and quantitative evidence that the Laplace mechanism is preferable to approximate DP for privatizing per-class per-leaf counts in this setting.
III Reconstruction Attack against DP RFs
@@
Adversary model. To the best of our knowledge, our attack is the first method to target full dataset reconstruction against DP-protected models. We assume white-box access to an $\varepsilon$-DP RF constructed using the algorithm of [24], as described in Section II-B. This includes access to the structure of the forest (split attributes and values) and the (noisy) counts at the leaves. Consistent with the literature on reconstruction attacks [5, 9], we also assume knowledge of the number of attributes $M$ and training examples $N$. While the knowledge of $M$ is necessary for inference, the knowledge of $N$ simplifies the attack but is not strictly required. We briefly discuss how the proposed CP formulation can be adapted to handle this setting at the end of this section, and provide the detailed modified formulation in Appendix A. We also assume the privacy budget $\varepsilon$ is known. This does not degrade the DP guarantees, is generally disclosed by practitioners [^2] For instance, [25] documents real-world applications of DP and their associated budgets., and researchers have advocated this for transparency [26].
@@
Reconstruction approach. Our approach is grounded on the CP formulation introduced in [9], with additional constraints, variables and an objective designed to model the (observed) noisy counts, the underlying (guessed) true counts and their likelihood. Solving the original problem formulation (DRAFT) was shown to be NP-Hard. Our reconstruction problem for DP RFs can be seen as a generalization (i.e., DRAFT is a special case with no noise) and is also NP-hard, though scalable with state-of-the-art CP solvers [27].
@@
Our formulation revolves around three main groups of decision variables. The objective of a CP solver is to find compatible variable values that satisfy a given set of constraints, yielding a feasible solution. Moreover, as explained later in this section, an objective function steering the search towards likely solutions (reconstructions) will be included. First, for each tree $t\in\mathcal{T}$, the (guessed) true number of examples of class $c\in\mathcal{\mathcal{C}}$ in each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are encoded using a variable $n_{tvc}\in\mathbb{N}$. Next, a set of variables encodes each reconstructed example $k\in\{1,\dots,N\}$: $x_{ki}\in\{0,1\}$ encodes the value of its $i^{\text{th}}$ attribute $i\in\{1,\dots,M\}$, while $z_{kc}\in\{0,1\}$ indicates if it belongs to class $c\in\mathcal{C}$. Third, the path of each example $k\in\{1,\dots,N\}$ through each tree $t\in\mathcal{T}$ is modelled using a binary variable $y_{tvkc}\in\{0,1\}$, which is $1$ if it is classified by leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ as class $c\in\mathcal{C}$, and 0 otherwise. The following constraints linking these variables are then enforced:
$\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}n_{tvc}=N$ $t\in\mathcal{T}$ (1)
$\sum\limits_{c\in\mathcal{C}}z_{kc}=1$ $k\in\{1,\dots,N\}$ (2)
$z_{kc}=\sum\limits_{\smash{\mathclap{v\in\smash{\mathcal{V}_{t}^{L}}}}}y_{tvkc}$ $k\in\{1,\dots,N\},t\in\mathcal{T},c\in\mathcal{C}$ (3)
$\sum_{k=1}^{\smash{N}}y_{tvkc}=n_{tvc}$ $t\in\mathcal{T},~v\in\smash{\mathcal{V}_{t}^{L}},~c\in\mathcal{C}.$ (4)
@@
In a nutshell, Constraint (1) ensures that the example counts sum up to $N$ within each tree. Constraint (2) guarantees that each example is associated to exactly one class. Constraint (3) enforces that each example contributes only to the counts of its assigned class. Moreover, Constraint (4) ensures consistency between the counts at each leaf and the examples assigned to it. Finally, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ of each tree $t\in\mathcal{T}$, let $\Phi_{v}^{+}$ (respectively $\Phi_{v}^{-}$) be the set of binary attributes that must equal $1$ (respectively 0) for an example to fall into $v$. Then, the following constraint enforces consistency between the examples’ assignments to the leaves and the value of their attributes: For $t\in\mathcal{T},~v\in\smash{\mathcal{V}_{t}^{L}},~k\in\{1,\dots,N\},$
@@
$\sum\limits_{c\in\mathcal{C}}y_{tvkc}=1\Rightarrow$ $\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{+}}x_{ki}=1\right)\wedge\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{-}}x_{ki}=0\right).$ (5)
@@
$p_{l}=\mathbb{P}(\textrm{int}(Y_{tvc})=l)=\begin{cases}\int_{l}^{l+1}\frac{\varepsilon_{v}}{2}e^{-\lvert u\rvert\varepsilon_{v}}\,du&\text{if }l>0,\\ \int_{-1}^{1}\frac{\varepsilon_{v}}{2}e^{-\lvert u\rvert\varepsilon_{v}}\,du&\text{if }l=0,\\ \int_{l-1}^{l}\frac{\varepsilon_{v}}{2}e^{-\lvert u\rvert\varepsilon_{v}}\,du&\text{if }l<0.\end{cases}$
In practice, we pre-compute constants $p_{l}$ for all values $l\in\{-\gamma,\dots,\gamma\}$ such that $\mathbb{P}(\textrm{int}(Y_{tvc})\in\{-\gamma,\dots,\gamma\})\geq 0.999$. Details on the computation of $\gamma$ are provided in Appendix B, but typically, $\gamma=\lceil 12/\varepsilon_{v}\rceil$ is sufficient. With this, we can additionally restrict the feasible domain of the variables to $n_{tvc}\in\{\max(0,n_{tvc}^{*}-\gamma),\dots,n_{tvc}^{*}+\gamma\}$ and $\Delta_{tvc}\in\{-\gamma,\dots,\gamma\}$ and formulate the objective as:
$\max\ \sum\limits_{t\in\mathcal{T}}\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}\sum\limits_{l=-\gamma}^{\gamma}\log(p_{l})\mathds{1}_{\Delta_{tvc}=l},$ (9)
@@
Reconstruction approach when $N$ is unknown. While the CP formulation introduced earlier assumes knowledge of the number of training examples $N$, it can actually be adapted to the more challenging case where $N$ is not known. Appendix A provides the full modified formulation along with a detailed empirical evaluation. The high-level idea is as follows. First, we leverage the knowledge of the distribution of the noise added to each individual leaf count to estimate the variance of their sum for each individual tree (whose true value without noise should be the unknown $N$). We consequently use the Central Limit Theorem to determine a 9595% confidence interval $\smash{\llbracket N_{\textsc{min}},N_{\textsc{max}}\rrbracket}$ for $N$, given the average sum of the noisy counts for the different trees in the forest. In the modified CP formulation, we allow up to $N_{\textsc{max}}$ examples to be reconstructed, with the actual number determined by an additional integer variable $\smash{\widetilde{N}}\in{N_{\textsc{min}},\dots,N_{\textsc{max}}}$. In the encoded reconstruction, only examples $k$ such that $\smash{k\leq\widetilde{N}}$ contribute to the inferred leaf counts. The objective function (9) then drives $\smash{\widetilde{N}}$ towards the true $N$ by maximizing the likelihood of the inferred noise values.
Finally, these modifications increase the size of the CP formulation search space, making the problem computationally more challenging. Nevertheless, our empirical evaluation shows that it can still be solved efficiently and achieves reconstruction errors comparable to those obtained when $N$ is known.
IV Experimental Study